OpenStudy (anonymous):

write a quadratic equation in the variable x having the given numbers as a solution. ax^2+bx+c=0 2, only solution

OpenStudy (anonymous):

(x - 2)(x - 2) Can you FOIL it to get the answer?

OpenStudy (anonymous):

if x=2 is the only solution of ax^2+bx+c=0, then a(x-2)^2=0, a=any number except 0 or 2*b+c=0, a=0 c=-2b. in the first case, a(x-2)^2=0 → ax^2-4ax+4a=0 → b=-4a, c=4a

OpenStudy (anonymous):

x^2-4x+4=0